Solve the equation by completing the square
x^2 - 14x + 24 = 0

Answer by ankor@dixie-net.com(22740)

Solve the equation by completing the square
x^2 - 14x + 24 = 0
:
x^2 - 14x + ___ = -24
:
Find the 3rd term to make it a perfect square. Divide the coefficient of x by 2 and square it:
In this problem: 14/2 = 7; 7^2 = 49, (add 49 to both sides)
x^2 - 14x + 49 = -24 + 49
:
In factored form we have:
(x - 7)^2 = 25
:
Find the square root of both sides:
x - 7 = +/-Sqrt(25)
:
x - 7 = +/-5
:
Two solutions:
x = + 7 + 5
x = 12
and
x = +7 -5
x = 2
:
:
Check our solutions by FOILing (x-12)*(x-2) = x^2 - 14x + 24